OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)?

OpenStudy (anonymous):

1st find the distance of centre using distance formula... then use this (x-x1)^2+(y-y1)^2=c^2 c=centre x1=-2 y1=0

OpenStudy (anonymous):

(x − 2)2 + (y + 3)2 = 5 (x + 2)2 + (y − 3)2 = 25 (x − 2)2 + (y + 3)2 = 25 (x − 2)2 − (y + 3)2 = 5

OpenStudy (anonymous):

those are the answer options

OpenStudy (jdoe0001):

 use the distnace formula to find the radius firstly, thus \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ center&({\color{red}{ 2}}\quad ,&{\color{blue}{ -3}})\quad point&({\color{red}{ -2}}\quad ,&{\color{blue}{ 0}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\) keep in mind that \(\bf (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2={\color{olive}{ r}}^2\qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad radius={\color{olive}{ r}}\)

OpenStudy (jdoe0001):

hmmm so \(\b \bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ center&({\color{red}{ 2}}\quad ,&{\color{blue}{ -3}})\quad point&({\color{red}{ -2}}\quad ,&{\color{blue}{ 0}}) \end{array} \\ \quad \\ d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

I'm so confused. would it be a=5 B=25 C=25 or D=5

OpenStudy (jdoe0001):

?

OpenStudy (anonymous):

A (x − 2)2 + (y + 3)2 = 5 B (x + 2)2 + (y − 3)2 = 25 C (x − 2)2 + (y + 3)2 = 25 D (x − 2)2 − (y + 3)2 = 5

OpenStudy (jdoe0001):

so... what did you get for the radius?

OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)

OpenStudy (jdoe0001):

we read that, yes

OpenStudy (anonymous):

thats the question it isn't asking for a radius, it wants the standard equation.. thats why I said A B C or D

OpenStudy (anonymous):

it was C i guessed :)

OpenStudy (anonymous):

A circle is described by the equation x2 + y2 + 14x + 2y + 14 = 0. What are the coordinates for the center of the circle and the length of the radius?

OpenStudy (anonymous):

(-7, -1), 36 units (7, 1), 36 units (7, 1), 6 units (-7, -1), 6 units